OpenStudy (mason):

Angela ate at a restaurant four times. Each time she ordered a salad and left a $5 tip. She spent a total of $54. Write and solve an equation to find the cost of each salad.

OpenStudy (phi):

what do you have to "solve for" ?

OpenStudy (mason):

solve for the cost of the salads

OpenStudy (phi):

ok, then in algebra, we give what we are solving for a short "name" usually "x" (for unknown)

OpenStudy (mason):

ya

OpenStudy (mason):

so solve for x salads. But how do i do

OpenStudy (phi):

if one salad costs x any idea how much 4 salads will cost ?

OpenStudy (mason):

ummm idk

OpenStudy (phi):

it's a way of thinking. if one salad costs 3 (dollars) how much does 4 salads cost ? any idea ?

OpenStudy (mason):

ya but what if it isnt 3 bucks

OpenStudy (mason):

it dosnt say

OpenStudy (phi):

just answer the question: if one salad costs 3 (dollars) how much does 4 salads cost ? any idea ?

OpenStudy (mason):

12

OpenStudy (phi):

now, the thinking part. How did you get 12 ?

OpenStudy (mason):

3*4

OpenStudy (phi):

right, and if the cost is x (instead of 3) we *use the same idea* how much do 4 salads cost if one salad costs x ?

OpenStudy (mason):

3

OpenStudy (mason):

i think

OpenStudy (phi):

how much does 4 salads cost if one salad costs 3 you said 4*3 now use that same idea for how much does 4 salads cost if one salad costs x (and remember, you are allowed to multiply a letter)

OpenStudy (mason):

4*x

OpenStudy (phi):

yes, or 4x (we usually leave out the multiply sign,but 4x means 4 * x)

OpenStudy (mason):

aight

OpenStudy (phi):

now the cost of the meal also included a 5 tip in other words, the cost of one meal was the cost of the salad x plus the tip 5 can you write that in algebra ?

OpenStudy (mason):

4x+5

OpenStudy (mason):

correct?

OpenStudy (phi):

you wrote 4 salads and 1 tip but I think she left a tip at each meal so that is close (but not quite right)

OpenStudy (mason):

4x+20

OpenStudy (mason):

correct?

OpenStudy (phi):

yes, and that equals what ? (they tell you in the question)

OpenStudy (mason):

54

OpenStudy (phi):

yes, so you can write an equation showing one side equals the other

OpenStudy (mason):

what is the cost of each salad tho?

OpenStudy (phi):

they ask *** Write and solve an equation **** we are doing the first part, writing down the equation

OpenStudy (mason):

so i wrote 4x+20, now what

OpenStudy (phi):

that is not the whole equation (which must have an = sign) in other words, it "should make sense" that 4x+20 = what ??

OpenStudy (mason):

54

OpenStudy (phi):

yes, so please write down the equation (which has an equal sign in it)

OpenStudy (mason):

4x+20=54

OpenStudy (mason):

correct

OpenStudy (phi):

yes, that looks good. Try to always write down the whole equation. now we solve it. first step is to "add" -20 to both sides, like this 4x+20 -20 = 54-20 now "simplify" both sides. can you do that ?

OpenStudy (mason):

no

OpenStudy (phi):

on the right side you have 54-20 you can simplify that ?

OpenStudy (mason):

34

OpenStudy (phi):

and on the left side you have a 20 -20 what is that ?

OpenStudy (mason):

0

OpenStudy (phi):

ok, so what does the equation look like now? try to write it down

OpenStudy (mason):

34(something)10

OpenStudy (phi):

4x+20 -20 = 54-20 try again.

OpenStudy (mason):

4x=34

OpenStudy (phi):

yes (I assume you know 4x+0 is just 4x)

OpenStudy (mason):

so thats right??

OpenStudy (phi):

that equation says: 4 salads cost 34 (dollars) but using algebra, we divide both sides by 4 can you try doing that?

OpenStudy (phi):

\[ \frac{4x}{4} = \frac{34}{4} \]

OpenStudy (mason):

ya

OpenStudy (phi):

on the left side 4/4 is 1 so you get 1x which we usually write a just x what do you get on the right side ?

OpenStudy (mason):

idk

OpenStudy (phi):

\[ \frac{4x}{4} = \frac{34}{4} \] x= 34/4 you have to figure out 34/4

OpenStudy (mason):

8.4

OpenStudy (phi):

you should write the whole equation x= 8.40 (add the zero, because this in money) 8 dollars and 40 cents.

OpenStudy (mason):

so what would be the equation

OpenStudy (phi):

we figured it out up above. It is what you solved.

OpenStudy (mason):

4x+20

OpenStudy (phi):

almost. but that is not an equation remember it has to have an = sign and another side

OpenStudy (mason):

4x+20=54

OpenStudy (phi):

yes

OpenStudy (mason):

so thats my answer

OpenStudy (phi):

yes 4x+20=54 is the equation after you solve it you get x= 8.40

OpenStudy (mason):

thx dude, and do you do this for money???

OpenStudy (phi):

no, an engineer who uses this stuff

OpenStudy (mason):

thx anyways dude

OpenStudy (mason):

cya :3